Project

General

Profile

Meeting #16629

Meeting notes (running)

Added by Jonathan Edelen over 3 years ago. Updated almost 3 years ago.

Status:
New
Priority:
Normal
Assignee:
-
Start date:
05/23/2017
Due date:
% Done:

0%

Estimated time:
Duration:

History

#1 Updated by Jonathan Edelen over 3 years ago

Meeting 22 May 2017

  • We had a meeting about the MEBT chopper with Greg, Daniil, Ed, Jonathan, and Brian.
  • Discussed clock issues (see below for Brian's email
    • This is a suggestion that we keep the controls clock frequency at 9.0277... MHz (162.5 MHz/18)
  • There is a plan to change the clock system to 10 MHz from the current 9.027 MHz provided by LLRF. 10 MHz has many advantages to users, however, yesterday in a MEBET kicker meeting, it became clear that going away from a subharmonic of the 162.5 MHz beam frequency will make it impossible to provide a distributed event and trigger system that will stay aligned to the pulsed beam pattern produced by the chopper system. Losing this relationship with the beam pattern will challenge beam instrumentation systems, LLRF beam loading compensation, Beam Chopper learning, Booster injection kickers timing and any future beam experiments.
  • For PIP-II itself it would be valuable to increase the clock frequency to a higher RF based harmonic with newer electronics. Now, it is important to maintain the harmonic relationship with the beam. This may warrant more discussion.
  • Action items
    • LLRF will provide a trigger signal (may be from controls)
    • Need to study the phase alignment of the clock on re-boot of the system
    • Greg and Co will have cables pulled to rack 200
    • Plan is to do initial testing using tectronix
    • Jonathan work on the block diagram
  • Keep the 50 Ohm kicker in the back of our mind

#2 Updated by Jonathan Edelen about 3 years ago

Meeting notes 5 July 2017

  • 200 Ohm kicker and 50 Ohm kicker are installed and we have all the permissions to run
  • Initial testing using DC voltages on 200 Ohm kicker, ~ 16% lower kick than expected
    • Measurements are ongoing for qualification
  • Working on three major components, synchronization circuit, comparator circuit, LabVIEW application to program the ARB
  • Status of comparator circuit
    • Circuit is finished signal levels seem good
  • Status of synchronization circuit
    • Tested with one-shot, close to completion
    • Should be ready by the end of today (5 July)
  • Status of software
    • Also close to completion
    • There are some issues with the filter, need to lower the gain to avoid saturation
  • Need a computer for the labview application that can be controlled remotely from the control room

#3 Updated by Joshua Einstein almost 3 years ago

Meeting Notes 25 October 2017

  • LV application for pattern generation has been working well. 50 Ohm in development.
  • All necessary hardware is now available. Possibly missing a synchronization box necessary for trigger-to-RF synchronization.

#4 Updated by Joshua Einstein almost 3 years ago

Meetign Notes 10 November 2017

  • Meeting for IIFC collaboration tomorrow
    • See IIFC collaboration project for notes from the meeting


Also available in: Atom PDF